# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v2

@@ 4.4.3. Provenance verification process
The first step is to derive an inference threshold $c$ based on a pre-defined FPR ($\alpha$). $\alpha$ is a small number such as 0.1% or 0% to ensure low false positive (e.g., Section 5 shows that MembershipTracker can support the users to achieve 100% TPR with 0% FPR in many cases). To derive the threshold $c$, $U$ first computes the empirical loss histogram for data in the out world $\mathcal{P}_{\text{out}}$. For each non-member user in $\mathcal{P}_{\text{out}}$, the auditing user ${U}$ first computes the per-instance loss on their data (line 2), and then derives the average loss for each (line 3). Line 4 estimates the CDF for the loss distribution, and computes its $\alpha$-percentile to derive a threshold $c$ (Ye et al., 2022).
@@ 5. Evaluation
Datasets and model training. We consider six common benchmark datasets, including CIFAR100 (Krizhevsky et al., 2009), CIFAR10 (Krizhevsky et al., 2009), ArtBench (Liao et al., 2022), CelebA (Liu et al., 2015), TinyImageNet (Le and Yang, 2015) and ImageNet-1k (Deng et al., 2009). This section considers the first five datasets and the evaluation on the ImageNet-1k dataset is in Section 6. We follow standard model training practice such as using data augmentations and validation set - the model training details are in Appendix A.
@@
MembershipTracker configuration. We explain the default configurations for the main experiments, and we study different configurations in the ablation study (Appendix C).
@@
As explained earlier, we create outlier features as random color stripes, each with 16 stripes from 11 different common colors. Such a large feature set ($11^{16}$ choices) can support multiple users to generate their own features for data marking, and we use a image blending ratio $m=0.7$. Next, we follow Co et al. (Co et al., 2019) to generate random perlin noise for each sample with $\delta=8/255$. We provide additional visualizations of the original and marked samples in Appendix A.
We assume each user possesses a small number of samples from a given class, which constitute 0.1% of the training set (the detailed dataset sizes are in Appendix A).
@@
As explained in Section 3.1.2, we adopt FPR@100% TPR as the evaluation metric, and we use 5,000 non-member users for computing FPR. We use all the samples that are not used for training as the non-member set. Due to the limited size of the non-member set, the non-member users’ samples are drawn randomly from the non-member set (with replacement), and these are similarly marked with random outlier features and perlin noise.
@@
To perform MI, we directly use the cross-entropy loss as the signal function value, which is easy to compute by the users. We consider the global-threshold method by Carlini et al. (Carlini et al., 2022) as the baseline method; and we do not choose the shadow-model-based approaches (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022). The reason is that, even though these methods can achieve better results, they still yield limited increase of MI success (Section 4.1); and more importantly, training shadow models is prohibitively challenging for ordinary personal users to carry out (Section 3.1.1).
@@
The closest work to ours in improving the MI success without shadow-model calibration is Tramer et al. (Tramèr et al., 2022). However, their technique assumes the users can manipulate the data labels, which excludes the real-world scenarios where the data labels are assigned by the model creator or an external service (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Though MembershipTracker does not impose such an assumption, for completeness, we still quantitatively compare it with the method by Tramer et al. (Appendix E).
@@ 5.1. Results
Dataset FPR@100% TPR Model accuracy
CIFAR100 0.00 66.34 (-0.41)
CIFAR10 0.00 90.64 (-0.50)
CelebA 0.00 69.41 (-0.63)
ArtBench 0.00 59.46 (-0.42)
TinyImageNet 0.00 67.22 (-0.40)
As shown, with MembershipTracker, the users can reliably trace the provenance of their marked data with 0% FPR@100% TPR. In comparison, without MembershipTracker, performing standard instance-based MI on the unmodified target samples incurs much higher FPR, with an average of 47.52% FPR@100% TPR (we also validate that it similarly has low performance under the low FPR regime, e.g., $<1\%$ TPR@0.1% FPR). This is because: 1) the unmarked samples cannot be strongly memorized by the model; and 2) the instance-based MI process fails to leverage the collective information across the target samples by each user. These two limitations are overcome by MembershipTracker’s data marking and set-based MI process; and the detailed ablation study on these two components is presented in Appendix C.
@@
Evaluation on different architectures and training-set sizes. We next consider additional evaluation (using CIFAR100 dataset) on different model architectures (including WideResNet (Zagoruyko and Komodakis, 2016), ResNet (He et al., 2016), ResNext (Xie et al., 2017), SeNet (Hu et al., 2018), GoogleNet (Szegedy et al., 2015) and DenseNet (Huang et al., 2017)), as well as different training-set sizes (from 5,000 to 25,000). We observe a similar trend as before (thus we omit the discussion), and the detailed results are in Appendix B.
@@
Supporting multiple users requires the model to memorize more target samples (from different users). However, DL models are known to be capable of memorizing a large corpus of data (e.g., to encode sensitive information (Song et al., 2017; Chen and Pattabiraman, 2024a)). Likewise, in our case, the model still strongly memorizes the marked samples and enables MembershipTracker to achieve 0% FPR@100% TPR even under multiple target users.
@@ 5.2.2. Adversarial augmentation
We consider employing additional augmentation techniques during training to mitigate the model’s memorization on the marked samples. We study three strategies: (1) injecting Gaussian noise (we inject zero-mean noise with various standard deviation values (0.02$\sim$0.5)); (2) using Mixup (Zhang et al., 2018) (we use seven different alpha parameters from 0.1 to 1.5); and (3) using RandAugment (Cubuk et al., 2020) to perform multiple random data augmentations (we consider 3$\sim$15 different augmentations).
@@
Figure 6. Adv. augmentation by Gaussian noise injection and RandAugment (Cubuk et al., 2020). MembershipTracker is resilient to moderate input augmentation, and aggressive augmentation can increase the TPR, but also jeopardize model accuracy.
@@ 5.2.3. Model fine-tuning
We now evaluate how model fine-tuning can mitigate the model’s memorization on the target samples. We assume an ideal defender with access to a small set (equivalent to 10% of training set) of clean and unmodified samples. We fine-tune the model with various fine-tuning rates from 0.1$\sim$0.00001, with 20 epochs for each. The results are on the left of Fig. 7. When the fine-tuning rates are small ($<$0.05), the models still exhibit strong memorization on the target samples, and thus MembershipTracker maintains 0% FPR@100% TPR.
Using a larger fine-tuning rate (e.g., 0.05, 0.1) can update the model in a more aggressive manner and hence alleviate the model’s memorization on the target samples. This, however, also hurts the model’s accuracy. For a fine-tuning rate of 0.05, the model accuracy is reduced by 2.9%, and the FPR by MembershipTracker increases from 0% to 0.5%. A larger fine-tuning rate of 0.1 can further increase the FPR to 4.4%, but it also causes a 5.97% accuracy drop.
Figure 7. Defense evaluation on model fine-tuning and model pruning. MembershipTracker is resilient to moderate model fine-tuning and pruning. Aggressive countermeasures can slightly increase the FPR incurred by MembershipTracker, but at the cost of major accuracy degradation.
@@ 5.2.4. Model pruning
Another related countermeasure we consider is to prune the model’s parameters, with the goal of removing the parameters that are associated with the target samples. We use the popular pruning strategy by Han et al. (Han et al., 2015), which sets the parameters with small absolute values to zero and has minimal impact to the model’s performance. We evaluate different pruning ratios and the results are on the right of Fig. 7.
@@ 5.2.5. Model-level defenses under multi-target setting
Among different defenses, we find that model fine-tuning stands out as the most potent solution that outperforms other methods in terms of high mitigation performance with low accuracy drop.
Compared with other methods (e.g., MI defenses (Abadi et al., 2016; Tang et al., 2022; Chen and Pattabiraman, 2024b)), fine-tuning only needs to perform standard fine-tuning on a small set of samples, and it can increase the FPR@100% TPR to 21.4% (with a fine-tuning rate of 0.05) with a mere 2.09% accuracy drop.
@@ 5.2.6. Input detection
Naturally, reducing the marking intensity would attenuate the model’s memorization on the marked samples, and increase their prediction loss. But we find that the prediction loss on the member samples are still much lower than that on non-members, and thus MembershipTracker can overcome it via the set-based MI process (similar to Section 5.2.1), and maintain a 0% FPR@100% TPR [^4] However, there are certain scenarios where the reduced data-marking intensity would lead to lower auditing performance, and an evaluation for this is in Appendix C.1.2..
@@
Specifically, we train the detector on samples marked with random color stripes, and evaluate it on samples marked with the features from an OOD (TinyImageNet) dataset (this is an effective alternative strategy we study in Appendix C.1.3). The result is in the orange line in Fig. 9, and the detection performance degrades considerably under such a realistic setting (with only 0.66 AUC).
@@ 5.2.7. Output perturbation
6. Evaluation on ImageNet Training
@@
We find that even under the large-scale training setting, the marked samples can still be strongly memorized by the models, which enables MembershipTracker to achieve 0% FPR@100% TPR.
@@
MembershipTracker can still maintain 0% FPR@100% TPR, when the marking ratio is reduced from 0.01% to 0.005%. However, when the marking ratio is further reduced to 0.0025%, MembershipTracker experiences lower auditing performance - this trend is similar to that in our ablation study in Appendix C.1.1 and thus we defer the detailed discussion to Appendix C.1.1 (due to space constraints). With that said, even under such a challenging case, MembershipTracker still achieves reasonably good performance as shown in the blue line of Fig. 11. To the best of our knowledge, MembershipTracker is the first technique that can scale to support reliable data auditing in large-scale training setting.
Figure 11. MembershipTracker under full-sized ImageNet training with different portions of samples for data marking. Aggressively reducing the portion of samples marked by each user can lead to reduced performance, though MembershipTracker can still maintain reasonably high auditing performance even when each user can only mark merely 0.0025% of the training data (the blue line).
@@ 7. Discussion
To understand how this may affect MembershipTracker’s performance, we study a hypothetical scenario where all the non-member samples (used for data auditing) have been used for training the target model. Even under such a challenging case, we find that MembershipTracker can still maintain reliable auditing performance with 0% FPR@100% TPR. We explain the reason below.
Recall that MembershipTracker performs data marking on non-members using random outlier features. These random features are highly unlikely to be used for training (e.g., among $10^{16}$ possible random color stripe features), and thus they significantly increase the prediction loss of non-member samples, e.g., reducing their accuracy from $>$99% (this is high because these non-member samples are used for training) to $<$30% after data marking. Consequently, the marked non-member samples still exhibit much higher prediction loss than marked member samples, from which MembershipTracker still maintains 0% FPR@100% TPR for reliable data auditing.
@@ 7.2. Limitations
To understand how this affects MembershipTracker, we perform an experiment where we assume the model owner collects 50 samples from each user (0.2% of the training set), and he/she includes only a subset of samples from each user for training. Table 5 shows the results, where we consider five target users (we observe a similar trend when considering the support for more target users). The FPR incurred by MembershipTracker grows as the number of marked samples used for training decreases. This is because those samples that are not used for training would yield higher prediction loss, which in turn increases the average loss on the entire set of marked samples, and affects MembershipTracker. On the other hand, while using a subset of user data can degrade MembershipTracker’s performance, it also results in lesser data for model training and causes an accuracy drop (third column in Table 5).
@@
Despite our best effort in keeping the targeted changes from severely distorting the marked data, the artifacts created by MembershipTracker are still visible under close inspection (more visual examples are shown in Appendix A). However, performing manual inspection to filter out the specially-marked samples by MembershipTracker can still be a challenge, particularly under the large-scale training setting (e.g., with millions of images). With that said, creating more subtle targeted changes still remains an important venue for future investigation.

### v3

@@ 4.2.3. Provenance verification process
The first step is to derive an inference threshold $c$ based on a pre-defined FPR ($\alpha$). $\alpha$ is a small number such as 1% or 0% to ensure low false positive. To derive the threshold $c$, $U$ first computes the empirical loss histogram for data in the out world $\mathcal{P}_{\text{out}}$. For each non-member user in $\mathcal{P}_{\text{out}}$, the auditing user ${U}$ first computes the per-instance loss on their data (line 2), and then derives the average loss for each (line 3). Line 4 estimates the CDF for the loss distribution, and computes its $\alpha$-percentile to derive a threshold $c$ (Ye et al., 2022).
@@ 5. Evaluation
Datasets and model training. We consider six common benchmark datasets, including CIFAR100 (Krizhevsky et al., 2009), CIFAR10 (Krizhevsky et al., 2009), ArtBench (Liao et al., 2022), CelebA (Liu et al., 2015), TinyImageNet (Le and Yang, 2015) and ImageNet-1k (Deng et al., 2009). We first consider the first five datasets and the evaluation on the ImageNet-1k dataset is in Section 5.3. For CIFAR10 and CIFAR100, we train a WideResNet-28-4 model on 25,000 samples for 100 epochs, with a learning rate of 0.1 (decayed by 5 at epoch 60, 80, 90), a weight decay of 5e-4 and momentum of 0.9. For TinyImageNet and ArtBench, we fine-tune an ImageNet-pretrained ResNet-18 model on 25,000 samples. For CelebA, we first follow (cel, [n. d.]) to create a 307-class subset from the dataset, and then similarly fine-tune an ImageNet-pretrained model on 2,000 samples. For fine-tuning, we use a small fine-tuning rate of 0.01 with 30 epochs and a momentum of 0.9. In all datasets except CelebA (where we find that fine-tuning on the entire set yields better accuracy), we use 20% of the training data as the validation set. We also use common data augmentations such as random cropping and horizontal flip.
Figure 5. Visualization of the original and marked samples.
@@
MembershipTracker configuration. We explain the default configurations for the main experiments, and we study different configurations in the ablation study (Appendix B).
@@
As explained earlier, we create outlier features as random color stripes, each with 16 stripes from 11 different common colors. Such a large feature set ($11^{16}$ choices) can support multiple users to generate their own features for data marking, and we use a image blending ratio $m=0.7$. Next, we follow Co et al. (Co et al., 2019) to generate random perlin noise for each sample with $\delta=8/255$. We assume each user possesses a small number of samples from a given class, which constitute 0.1% of the training set.
Fig. 5 provides additional visualizations of the original and marked samples, and we quantitatively evaluate the visual quality of the images in Appendix A.
@@
Unless otherwise stated [^4] We will use a different metric in Section 5.2 for reasons we will explain there., we follow the common practice to use TPR@low FPR for evaluation and we use a 0% FPR threshold (to minimize false accusations). We use 5,000 non-member users for controlling the FPR and these samples may be collected by sampling random data from the data distribution (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022). We use all the samples that are not used for training as the non-member set. Due to the limited size of the non-member set, the non-member users’ samples are drawn randomly from the non-member set (with replacement), and these are similarly marked with random outlier features and perlin noise.
@@
To perform MI, we directly use the cross-entropy loss as the signal function value, which is easy to compute by the users. We consider the global-threshold method by Carlini et al. (Carlini et al., 2022) as the baseline method; and we do not choose the shadow-model-based approaches (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022). The reason is that, even though these methods can achieve better results, they still yield limited increase of MI success (Section 4); and more importantly, training shadow models is prohibitively challenging for ordinary personal users to carry out (Section 3.1.1).
@@
The closest work to ours in improving the MI success without shadow-model calibration is Tramer et al. (Tramèr et al., 2022). However, their technique assumes the users can manipulate the data labels, which excludes the real-world scenarios where the data labels are assigned by the model creator or an external service (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Though MembershipTracker does not impose such an assumption, for completeness, we still quantitatively compare it with the method by Tramer et al. (Appendix D).
@@ 5.1. Results
Dataset TPR@0% FPR Model accuracy
CIFAR100 100.00 66.34 (-0.41)
CIFAR10 100.00 90.64 (-0.50)
CelebA 100.00 69.41 (-0.63)
ArtBench 100.00 59.46 (-0.42)
TinyImageNet 100.00 67.22 (-0.40)
As shown, with MembershipTracker, the users can reliably trace the provenance of their marked data with 100% TPR@0% FPR. In comparison, without MembershipTracker, performing standard instance-based MI on the unmodified target samples cannot reliably identify the member samples under a 0% FPR threshold and it has 0% TPR. This is because: 1) the unmarked samples cannot be strongly memorized by the model; and 2) the instance-based MI process fails to leverage the collective information across the target samples by each user. These two limitations are overcome by MembershipTracker’s data marking and set-based MI process; and the detailed ablation study on these two components is presented in Appendix B.
@@
Table 2. Evaluation on different models.
Model TPR@0% FPR Model accuracy
WideResNet 100.00 66.34 (-0.41)
ResNet 100.00 64.49 (-0.69)
ResNext 100.00 63.43 (-0.47)
DenseNet 100.00 67.14 (-0.16)
GoogleNet 100.00 69.09 (-0.33)
SeNet 100.00 65.26 (-0.37)
Table 3. Evaluation on different training-set sizes.
Training-set size TPR@0% FPR Model accuracy
25,000 100.00 66.34 (-0.41)
20,000 100.00 63.73 (-0.16)
15,000 100.00 58.99 (-0.83)
10,000 100.00 50.28 (-0.09)
5,000 100.00 35.74 (-0.10)
Evaluation on different architectures and training-set sizes. We next consider additional evaluation (using CIFAR100 dataset) on different model architectures (including WideResNet (Zagoruyko and Komodakis, 2016), ResNet (He et al., 2016), ResNext (Xie et al., 2017), SeNet (Hu et al., 2018), GoogleNet (Szegedy et al., 2015) and DenseNet (Huang et al., 2017)), as well as different training-set sizes (from 5,000 to 25,000). We report the results in Table 2 and Table 3, where we observe a similar trend as before (and thus we omit the discussion).
@@
Supporting multiple users requires the model to memorize more target samples (from different users). However, DL models are known to be capable of memorizing a large corpus of data (e.g., to encode sensitive information (Song et al., 2017; Chen and Pattabiraman, 2024a)). Likewise, in our case, the model still strongly memorizes the marked samples and enables MembershipTracker to achieve 100% TPR@0% FPR even under multiple target users.
@@ 5.2. Robustness to Countermeasures
Evaluation metric. In this section, we use a different evaluation metric, FPR@high TPR. This is because using TPR@low FPR (as in our previous evaluation) requires choosing a low FPR threshold and evaluating the corresponding TPR. However, the pre-defined FPR threshold may be too conservative or aggressive.
We use the results from the DPSGD experiment (Section 5.2.1) to explain the issue involved. In this example, using a 0.1% FPR threshold is too conservative as MembershipTracker has the same 100% TPR for the noise multiplier $\sigma=0.01$ and $\sigma=0.03$. Likewise, using a 0% FPR threshold is too aggressive as MembershipTracker has the same 0% TPR for $\sigma=0.2$ and $\sigma=0.3$. In both cases, using a pre-defined TPR may not adequately reflect MembershipTracker’s performance changes in different settings.
To overcome the above issue, we advocate the use of FPR under 100% TPR. This metric is able to properly compare MembershipTracker’s performance under different settings (e.g., Table 6), and hence we use it for the evaluation in this Section, as well as Section 6.2 and Appendix B (for the same reason explained above).
5.2.1. MI defense
@@ 5.2.2. Adversarial augmentation
We consider employing additional augmentation techniques during training to mitigate the model’s memorization on the marked samples. We study three strategies: (1) injecting Gaussian noise (we inject zero-mean noise with various standard deviation values (0.02$\sim$0.5)); (2) using Mixup (Zhang et al., 2018a) (we use seven different alpha parameters from 0.1 to 1.5); and (3) using RandAugment (Cubuk et al., 2020) to perform multiple random data augmentations (we consider 3$\sim$15 different augmentations).
@@
Figure 6. Adversarial augmentation by Gaussian noise injection and RandAugment (Cubuk et al., 2020).
@@
Figure 7. Evaluation with model fine-tuning. Fine-tuning the model with additional data can improve model accuracy (the more data the better), but MembershipTracker consistently maintains high auditing performance regardless.
@@ 5.2.3. Model fine-tuning
Fine-tuning is a technique performed on additional data that differs from the training data to improve model performance. We evaluate how it may be used to mitigate the model’s memorization on the target samples. We assume an ideal defender with access to an additional set of clean and unmodified samples. We consider multiple settings with different amounts of data (5,000 to 20,000 samples). For each setting, we fine-tune the model using various learning rates from 0.1$\sim$0.000001 for 20 epochs, and we report the one with the highest accuracy. The results are reported in Fig. 7.
As shown, using a small set of fine-tuning data (5,000 samples) can increase the model accuracy, and using more data can lead to larger accuracy improvement (understandably so). For instance, using 20,000 samples can increase the accuracy to 69.66%, which translates to a 10% reduction in the test error. Nevertheless, despite its effectiveness in enhancing model accuracy, model fine-tuning still has limited effectiveness in degrading MembershipTracker, which consistently achieves 0% FPR@100% TPR in all cases.
@@ 5.2.4. Model pruning
Another related countermeasure we consider is to prune the model’s parameters, with the goal of removing the parameters that are associated with the target samples. We use the popular pruning strategy by Han et al. (Han et al., 2015), which sets the parameters with small absolute values to zero and has minimal impact to the model’s performance. We evaluate different pruning ratios and the results are shown in Fig. 8.
@@
Figure 8. Evaluation with model pruning. MembershipTracker is resilient to moderate model pruning. Aggressive countermeasures can slightly increase the FPR incurred by MembershipTracker, but at the cost of major accuracy degradation.
@@ 5.2.5. Model-level defenses under multi-target setting
Among different defenses, we find that model fine-tuning stands out as the most potent solution that outperforms other methods in terms of high mitigation performance without accuracy drop. By fine-tuning the model on 20,000 additional samples, it can increase the FPR@100% TPR to 42.84% without accuracy drop (the model still has 67.30% accuracy).
@@ 5.2.6. Input detection
Naturally, reducing the marking intensity would attenuate the model’s memorization on the marked samples, and increase their prediction loss. But we find that the prediction loss on the member samples are still much lower than that on non-members, and thus MembershipTracker can overcome it via the set-based MI process (similar to Section 5.2.1), and maintain a 0% FPR@100% TPR [^5] However, there are certain scenarios where the reduced data-marking intensity would lead to lower auditing performance, and an evaluation for this is in Appendix B.1.2..
@@
Specifically, we train the detector on samples marked with random color stripes, and evaluate it on samples marked with the features from an OOD (TinyImageNet) dataset (this is an effective alternative strategy we study in Appendix B.1.3). The result is in the orange line in Fig. 10, and the detection performance degrades considerably under such a realistic setting (with only 0.66 AUC).
@@ 5.2.7. Output perturbation
5.3. Evaluation on ImageNet Training
@@
We find that even under the large-scale training setting, the marked samples can still be strongly memorized by the models, which enables MembershipTracker to achieve 100% TPR@0% FPR.
@@
MembershipTracker can still maintain 100% TPR@0% FPR, when the marking ratio is reduced from 0.01% to 0.005%. However, when the marking ratio is further reduced to 0.0025%, MembershipTracker experiences lower auditing performance - this trend is similar to that in our ablation study in Appendix B.1.1 and thus we defer the detailed discussion to Appendix B.1.1 (due to space constraints). With that said, even under such a challenging case, MembershipTracker still achieves reasonably good performance with $>$75% TPR@0.1% FPR (the blue line of Fig. 12). To the best of our knowledge, MembershipTracker is the first technique that can scale to support reliable data auditing in large-scale training setting.
Figure 12. Evaluating MembershipTracker under full-sized ImageNet training with different portions of samples for data marking.
@@ 6. Discussion
To understand how this may affect MembershipTracker’s performance, we study a hypothetical scenario where all the non-member samples (used for data auditing) have been used for training the target model. Even under such a challenging case, we find that MembershipTracker can still maintain reliable auditing performance with 100% TPR@0% FPR. We explain the reason below.
Recall that MembershipTracker performs data marking on non-members using random outlier features. These random features are highly unlikely to be used for training (e.g., among $10^{16}$ possible random color stripe features), and thus they significantly increase the prediction loss of non-member samples, e.g., reducing their accuracy from $>$99% (this is high because these non-member samples are used for training) to $<$30% after data marking. Consequently, the marked non-member samples still exhibit much higher prediction loss than marked member samples, from which MembershipTracker still maintains 100% TPR@0% FPR for reliable data auditing.
@@ 6.2. Limitations
To understand how this affects MembershipTracker, we perform an experiment where we assume the model owner collects 50 samples from each user (0.2% of the training set), and he/she includes only a subset of samples from each user for training. Table 7 shows the results, where we consider five target users (we observe a similar trend when considering the support for more target users). We use FPR@100% TPR as the metric, as explained earlier in Section 5.2.
The FPR incurred by MembershipTracker grows as the number of marked samples used for training decreases. This is because those samples that are not used for training would yield higher prediction loss, which in turn increases the average loss on the entire set of marked samples, and affects MembershipTracker. On the other hand, while using a subset of user data can degrade MembershipTracker’s performance, it also results in lesser data for model training and causes an accuracy drop (third column in Table 7).
@@
Despite our best effort in keeping the targeted changes from severely distorting the marked data, the artifacts created by MembershipTracker are still visible under close inspection (e.g., Fig. 5). However, performing manual inspection to filter out the specially-marked samples by MembershipTracker can still be a challenge, particularly under the large-scale training setting (e.g., with millions of images). With that said, creating more subtle targeted changes still remains an important venue for future investigation.